# Discontinuous integro-differential equations and sliding mode control

### v2

@@ 5 Conclusions
In this paper, Filippov method is extended to systems modeled by integro-differential equations. This method allows an ill-posed discontinuous system to be regularized by means of its transformation to a well-posed differential inclusion. Such ill-posed discontinuous systems appear in sliding mode control theory, which uses discontinuous feedback laws for regulation of systems and rejection of matched perturbations. The advantage of the Filippov method consists in a minimal (from set-theoretic point if view) extension of an ill-posed equation to a well-posed equation. This implies a better matching of the dynamics of the control system by the regularized model. In the paper, the equivalent control method is developed for affine-in-control systems modeled by integro-differential equations. The well-posedness of regularized systems is studied. Filippov method and equivalent control approach are the main tools for analysis of control systems with sliding modes.
@@
A SMC for a plant modeled by infinite-dimensional differential equation with finite-dimensional inputs and outputs can also be designed using the developed theory of discontinuous IDEs. Such a discontinuous infinite-dimensional control system can be regularized based on a finite-dimensional IDE representing its input-output dynamics. This scheme is demonstrated on a heat control system modeled by PDE.
The obtained theoretical results provide all necessary tools for extension of other methods of SMC theory (e.g., high order SMC [Levant2003:IJC] or implicit SMC discretization [AcaryBrogliato2010:SCL]) to systems modeled by integro-differential equations and by differential equations in Banach spaces (e.g., PDEs).

### v3

@@ 5 Conclusions
In this paper, Filippov method is extended to systems modeled by integro-differential equations. This method allows an ill-posed discontinuous system to be regularized by means of its transformation to a well-posed differential inclusion. Such ill-posed discontinuous systems appear in sliding mode control theory, which uses discontinuous feedback laws for regulation of systems and rejection of matched perturbations. The advantage of the Filippov method consists in a minimal (from set-theoretic point of view) extension of an ill-posed equation to a well-posed inclusion. This implies a better matching of the dynamics of the control system by the regularized model. In the paper, the equivalent control method is developed for affine-in-control systems modeled by integro-differential equations. The well-posedness of regularized systems is studied. Filippov method and equivalent control approach are key tools for analysis of control systems with sliding modes.
@@
For a class of infinite-dimensional control systems (64) with finite-dimensional inputs and outputs, a sliding mode control can be designed and analyzed using the developed theory. Indeed, the input-output dynamics of such an infinite-dimensional system can always be represented as a finite-dimensional integro-differential equation. If the input-output dynamics admits the representation (46), (47), then the SMC for the original system can be designed in the form (48) provided that all conditions of Theorem 20 are fulfilled. In Example 8, this approach is demonstrated for a heat control system modeled by PDE.
The obtained theoretical results provide all necessary tools for extension of other methods of SMC theory (e.g., high order SMC [29] or implicit SMC discretization [1]) to systems modeled by integro-differential equations and by differential equations in Banach spaces (e.g., PDEs).